OpenStudy (anonymous):

Solve: 1/y = 2/y-1/5 the solution is ?

OpenStudy (anonymous):

i got 5?????? is that right?????

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

y=-5/2

OpenStudy (anonymous):

You're incorrect, the inversion step is wrong:\[1/y=2/y-1/5\neq>y= y/2-5\]

OpenStudy (anonymous):

\[{1 \over y}={2 \over y-{1 \over 5}} \implies y-{1 \over 5}=2y \implies y=-{1 \over 5}\]

OpenStudy (anonymous):

I don't think that's the equation. From what she put, it should be:\[1/y=2/y-1/5\]Not:\[\frac{1}{y}=\frac{2}{y-1/5}\]

OpenStudy (anonymous):

You're right. But I am pretty sure she meant what I wrote, because she did the same thing in another post.

OpenStudy (anonymous):

Oh ok, it's just that the solution to the one I wrote is 5 so I thought it was probably that.

OpenStudy (anonymous):

Well, she got two solutions. It's her job now to pick the one that answers her question :)